# Controllable seismic velocity synthesis using generative diffusion models

### v1

@@ Abstract
Accurate seismic velocity estimations are vital to understanding Earth’s subsurface structures, assessing natural resources, and evaluating seismic hazards. Machine learning-based inversion algorithms have shown promising performance in regional (i.e., for exploration) and global velocity estimation, while their effectiveness hinges on access to large and diverse training datasets whose distributions generally cover the target solutions. Additionally, enhancing the precision and reliability of velocity estimation also requires incorporating prior information, e.g., geological classes, well logs, and subsurface structures, but current statistical or neural network-based methods are not flexible enough to handle such multi-modal information. To address both challenges, we propose to use conditional generative diffusion models for seismic velocity synthesis, in which we readily incorporate those priors. This approach enables the generation of seismic velocities that closely match the expected target distribution, offering datasets informed by both expert knowledge and measured data to support training for data-driven geophysical methods. We demonstrate the flexibility and effectiveness of our method through training diffusion models on the OpenFWI dataset under various conditions, including class labels, well logs, reflectivity images, as well as the combination of these priors. The performance of the approach under out-of-distribution conditions further underscores its generalization ability, showcasing its potential to provide tailored priors for velocity inverse problems and create specific training datasets for machine learning-based geophysical applications.
@@ 1 Introduction
Specifically designed datasets for machine learning-based velocity inversion, like the OpenFWI and OpenFWI2 datasets, have emerged recently. However, to scale the neural networks to accommodate real-world scenarios, the datasets often used possess several limitations: 1) They rely on manually defined velocities, which lack the necessary diversity for large-scale neural networks. 2) Synthetic datasets demand significant physical space for storage and sharing. In comparison to datasets explicitly tailored for the task, deep learning can be employed to generate velocities. Ovcharenko et al. (2019) proposed to use style transfer to generate velocities based on synthetic prior velocities. Alternatively, generative models, which have shown considerable potential in capturing the distribution of a dataset and enabling the generation of new samples, can be a more efficient data synthesis tool. Generative models have favorable advantages: 1) they can produce high-quality and diverse large datasets as their parameters are used to store the distribution; 2) they provide potentially unlimited training data from the stored distribution features; 3) they can utilize both OpenFWI datasets as well as other datasets (SEAM, Overthrust), even velocity models from industry, to generate more realistic velocities; 4) using the samples from the generative model for training can further improve the performance of NNs He et al. (2022). Puzyrev et al. (2022) proposed to use GANs to generate velocity models, but training GANs is fraught with difficulties, and their progress in generative tasks remains uncertain. In addition, the learned velocity distributions— leveraging these generative models— often lack essential foundational information regarding the velocity model classes (general geological structural description like salt, fault, flat, curve) or features, rendering them somewhat blind. In this paper, we use conditional diffusion models to synthesize seismic velocity models, which can provide desired and fruitful velocities, dependent on the target velocity classes, well logs, and structural information, as training datasets for data-driven geophysical methods.
@@
Beyond addressing the challenge of lacking training datasets whose distribution covers the target solution, another important challenge we often face is the inflexibility of incorporating multi-modal priors into FWI. Along these lines, Wang et al. (2023) demonstrated that injecting priors stored in diffusion models into the inversion process can improve the results. This is done by regularly updating the velocity model using diffusion denoising regularizer after each iteration, leading to increased accuracy and accuracy of the seismic velocity model. However, the priors used in their approach mainly provide high-level information, reflecting general velocity ranges and structures. While useful, this is often insufficient for achieving the desired level of accuracy in scenarios of poorly acquired data. In this case, additional prior information (geological classes, well logs, and structure information) is needed. Hence, the critical challenge lies in how to handle these multi-modal priors flexibly, and that can be addressed by controllable velocity synthesis.
@@
  • We propose to use conditional diffusion models for velocity synthesis to support downstream data-driven geophysical methods and diffusion-regularized FWI.
@@ 2 Methology
Generative models are widely used in seismic exploration, such as in denoising and interpolating seismic data Oliveira et al. (2018, 2019); Gatti and Clouteau (2020); Li et al. (2020); Feng et al. (2021); Wei et al. (2021, 2023), seismic inversion Mosser et al. (2020); Meng et al. (2021); Zhao et al. (2021); Wang et al. (2023), interpretation Shi et al. (2020); Durall et al. (2021b, a), and so on. These generative models mainly belong to four types of generative frameworks: variational autoencoders (VAEs) Kingma and Welling (2013), generative adversarial networks (GANs) Goodfellow et al. (2020), normalizing flows Dinh et al. (2014), and diffusion models Sohl-Dickstein et al. (2015). Among all generative models, VAE and normalizing flows require stringent constraints on the model architecture or resort to surrogate objectives for approximating the maximum likelihood training. Training GANs is fraught with difficulties, and their progress in generative tasks remains uncertain. Recently, diffusion models, which circumvent the above limitations, have beaten GANs in generative quality and diversity Dhariwal and Nichol (2021). Here, we choose diffusion models for seismic velocity synthesis.
The fundamental concept behind Denoising Diffusion Probabilistic Models (DDPMs) is straightforward: we aim to acquire the ability to restore images in the 2D context (or videos in the 3D realm) by training denoisers to reverse the corruption process, ultimately restoring the original data distribution from a Gaussian distribution of the latent space. This comprehensive framework can be delineated into two distinct stages: the training phase and the sampling phase, as depicted in Figure 1. In real-world applications, we aim to exert control over the velocity generation through conditional factors. This control can be introduced during the training process, which is referred to as ’classifier-free guidance’, or it can be implemented directly in the sampling stage, known as the ’classifier-guided (reconstruction-guided) diffusion model.’ In this section, we will initially delve into the training stage before exploring the sampling process in subsequent subsections.
@@ 2.1 Training process
During the training, given clean data $\boldsymbol{x}\sim p(\boldsymbol{x})$, we use a forward process (Gaussian process) to corrupt the data and obtain the noisy $\boldsymbol{x}_{t}$, $\boldsymbol{x}_{t}\sim p_{t}(\boldsymbol{x})$, which satisfies a linear stochastic differential equation (SDE):
@@
where $\beta(t)$ is a differentiable noise scheduler to enforce $\boldsymbol{x}_{1}$ to satisfy a pre-defined noise distribution (here it is the standard Gaussian distribution), $\boldsymbol{w}$ is the Wiener process (Brown motion, which adds random noise to data at each time step), and $d$ denotes the change in a variable, like $\boldsymbol{x}$, $t$, or $\boldsymbol{w}$ in Equation 1. As shown at the top of Figure 1, for a clean data sample $\boldsymbol{x}_{0}$, we gradually add noise to the data $\boldsymbol{x}_{t}$ until we obtain a final image $\boldsymbol{x}_{1}$ of random noise. The discrete form of Equation 1 (taking DDPM as an example) can be formulated as:
@@
$\mathbb{E}_{\boldsymbol{\epsilon},t}\left[\left(\frac{1+\bar{\alpha}_{i}}{1-\bar{\alpha}_{i}}\right)^{(-\gamma)}\left\|\Phi\left(\theta,\boldsymbol{x}_{i}\right)-\boldsymbol{x}\right\|_{2}^{2}\right],$ (3)
where $\gamma$ is the loss weight value Choi et al. (2022). In practice, we do not use the direct prediction of $\boldsymbol{x}$ via the neural networks. Instead, we use the $\boldsymbol{\epsilon}$-prediction parameterization, defined as $\Phi(\theta,\boldsymbol{x}_{i})=(\frac{\boldsymbol{x}_{i}}{\sqrt{\bar{\alpha}_{i}}}-\sqrt{\frac{1-\bar{\alpha}_{i}}{\bar{\alpha}_{i}}}\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i}))$ and train the $\boldsymbol{\epsilon}_{\theta}$ in $\boldsymbol{\epsilon}$ space. It results in the solution of a scaled score estimate $\boldsymbol{\epsilon}_{\theta}(z_{t})\approx-\sqrt{1-\bar{\alpha}_{t}}\nabla_{\boldsymbol{x}}\log p_{t}\left(\boldsymbol{x}\right)$.
@@ 2.2 Unconditional sampling process
To generate an unconditional sample, a diffusion model starts with a random Gaussian noisy image and then applies a sequence of denoising steps to reduce the noise gradually, as shown at the bottom of Figure 1. At each step, the model predicts a residual image that is added to the noisy image to produce a less noisy image. The process continues until the noise is sufficiently reduced and a high-quality sample is obtained. The sampling process can be described in the continuous form as solving a reverse-time SDE Song et al. (2020b)
@@ 2.3 Conditional sampling process
In the conditional sampling (conditional generation) setting, the generation process can be formulated in the context that given the conditioning signal $\boldsymbol{c}$, which can represent a class label, smoothness level, well information, imaging results, or any other type of condition. The model generates output $\boldsymbol{x}$ satisfying the distribution of $p(\boldsymbol{x})$ while consistent with the set condition. Two conditional samplers are introduced here, including the reconstruction-guided samplers and classifier-free guidance samplers. The former can be applied to any well-trained diffusion model, as the guidance only happens at the sampling stage, while the latter requires training the diffusion model to include the conditions.
Reconstruction guided generation. In this way, we assume that the unconditional generation sampler $\Phi(\theta,\boldsymbol{x})$ satisfies the conditions, in which we can use an objective function to guide the process:
@@
where $mask$ denotes the mask operation on the generated samples to extract, for our specific example, the well location information to compare with the condition given by the true well log. Then, the final samples at each time step $t$ is
$\Phi(\theta,\boldsymbol{x_{i}})\leftarrow\Phi(\theta,\boldsymbol{x}_{i})-\frac{w_{r}\sqrt{1-\alpha_{i}}}{2}\nabla_{\boldsymbol{x}_{i}}\left\|\boldsymbol{c}-mask(\Phi\left(\theta,\boldsymbol{x}_{i}\right))\right\|_{2}^{2},$ (7)
where the additional term with reweighting in this equation is the guidance based on the model’s reconstruction of the condition, so-called reconstruction-guided sampling, and $w_{r}$ is the weighting factor to control the strength of the condition. However, this process is very sensitive to the choice of the weighting factor, which we will demonstrate via the numerical results, yielding inflexible control.
Classifier-free guidance sampler. When considering the classes of the images, a concept similar to the reconstruction-guided diffusion model is called classifier guidance diffusion, whose generation process is guided by the gradient from an external classifier. Although it can feasibly guide the unconditional diffusion model with a plug-in classifier, this process will increase the computational cost as it requires gradient calculation during the inference. Another drawback of this type of method is that the generation sometimes will ignore the guidance of the classifier via the adversarial noise, then the generation in this case will actually be unconditional. Another controllable sampling is classifier-free guidance sampling. Compared to reconstruction-guided sampling, this requires training the diffusion model with the condition as part of the training. The good news, it often has higher-quality generation during the sampling stage. Our training target is to train a neural network to fit the distribution $p(\boldsymbol{x}|\boldsymbol{c})$ given a conditioning signal $\boldsymbol{c}$. Compared to the previously mentioned diffusion model (Figure 1), the only modification is changing the $\Phi(\theta,\boldsymbol{x})$ to $\Phi(\theta,\boldsymbol{x},\boldsymbol{c})$. To improve the sample quality, we do not directly use the prediction $\Phi(\theta,\boldsymbol{x},\boldsymbol{c})$. Instead, we use the following weighting form with the ability to adjust the strength of the conditioning,
@@
where $\lambda$ is the conditioning scale, and $\Phi(\theta,\boldsymbol{x}_{i},\boldsymbol{c})=(\frac{\boldsymbol{x}_{i}}{\sqrt{\bar{\alpha}_{i}}}-\sqrt{\frac{1-\bar{\alpha}_{i}}{\bar{\alpha}_{i}}}\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i},\boldsymbol{c}))$ is the conditional model prediction. When $\lambda$ is positive, the model output will have lower diversity, but higher quality compared to the case where the output is only dependent on the $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x},\boldsymbol{c})$. To make the diffusion model more flexible, in training, we add embedding vectors of the condition as well as the zero vector denoting unconditioning, yielding flexible generation abilities when sampling. Next, we will discuss how to incorporate the condition in the neural networks.
Figure 2: Conditional generative diffusion models.
As shown in Figure 2, there are two ways to achieve the classifier-free guidance. In the first way, the class label or the well log information is first embedded as a vector and then added into the hidden layer of the U-Net. However, this way has a limit in handling high-dimensional conditions, e.g., multiple wells or seismic images. In the second way, for 2D input data for conditioning, we utilize cross-attention to add conditional information into the neural networks for training. The concept is based on calculating the similarity between the conditioning data and the hidden output of the U-Net and then updating the latent vectors in the U-Net. We map the tokenizations of the original conditions $\mathbf{c}$, $\tau_{\theta}(\mathbf{c})$, to the intermediate output of the U-Net by implementing cross-attention$(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d}})\cdot V$, between the U-Net features and the input condition, where
@@
Here, $\phi_{j}(\boldsymbol{x}_{i})$ is the output of the intermediate layer of the U-Net, $\tau_{\theta}$ used in this paper is $1\times 1$ convolutional layer, and $W_{j}^{Q}$, $W_{j}^{K}$, and $W_{j}^{V}$ are the learnable weight matrices for the query, key, and value, respectively. To achieve an accurate control of velocity synthesis, we add the spatial information of the velocity (the coordinates of the grid points) as additional channels and then concatenate them with conditions $\boldsymbol{c}$ to feed into the $\tau_{\theta}$. The beauty of this method is that it can incorporate multi-model conditioning. What’s more, thanks to the cross-attention, we can achieve pixel-level controllable generation.

### v2

@@ Abstract
Accurate seismic velocity estimations are vital to understanding Earth’s subsurface structures, assessing natural resources, and evaluating seismic hazards. Machine learning-based inversion algorithms have shown promising performance in regional (i.e., for exploration) and global velocity estimation, while their effectiveness hinges on access to large and diverse training datasets whose distributions generally cover the target solutions. Additionally, enhancing the precision and reliability of velocity estimation also requires incorporating prior information, e.g., geological classes, well logs, and subsurface structures, but current statistical or neural network-based methods are not flexible enough to handle such multi-modal information. To address both challenges, we propose to use conditional generative diffusion models for seismic velocity synthesis, in which we readily incorporate those priors. This approach enables the generation of seismic velocities that closely match the expected target distribution, offering datasets informed by both expert knowledge and measured data to support training for data-driven geophysical methods. We demonstrate the flexibility and effectiveness of our method through training diffusion models on the OpenFWI dataset under various conditions, including class labels, well logs, reflectivity images, and the combination of these priors. The performance of the approach under out-of-distribution conditions further underscores its generalization ability, showcasing its potential to provide tailored priors for velocity inverse problems and create specific training datasets for machine learning-based geophysical applications.
Controllable seismic velocity synthesis using generative diffusion models
@@ 1 Introduction
Specifically designed datasets like the OpenFWI and OpenFWI2 datasets, as well as deterministic ways to build diverse velocity models Ren et al. (2021); Liu et al. (2021) for machine learning-based velocity inversion, have recently emerged. However, to scale the neural networks to accommodate real-world scenarios, the datasets often used possess several limitations: 1) They rely on manually defined velocities, which lack the necessary diversity for large-scale neural networks. 2) Synthetic datasets demand significant physical space for storage and sharing. In comparison to datasets explicitly tailored for the task, deep learning can be employed to generate velocities. Ovcharenko et al. (2019) proposed to use a style transfer approach to generate velocities based on synthetic prior velocities. Alternatively, generative models, which have shown considerable potential in capturing the distribution of a dataset and enabling the generation of new samples, can be a more efficient data synthesis tool. Generative models have favorable advantages: 1) they can produce high-quality and diverse large datasets as their learnable parameters are used to store the distribution; 2) they provide potentially unlimited training data from the stored distribution features; 3) they can utilize both OpenFWI datasets as well as other datasets (SEAM, Overthrust), even velocity models from industry, to generate more realistic velocities; 4) using the samples from the generative model for training can further improve the performance of NNs He et al. (2022). Puzyrev et al. (2022) proposed to use GANs to generate velocity models, but training GANs is fraught with difficulties, and their progress in generative tasks remains uncertain. In addition, the learned velocity distributions— leveraging these generative models— often lack essential foundational information regarding the velocity model classes (general geological structural description like salt, fault, flat, curve) or features, rendering them somewhat blind. In this paper, we use conditional diffusion models to synthesize seismic velocity models. It can provide desired and rich velocities, dependent on the target velocity classes, well logs, and structural information, as training datasets for data-driven geophysical methods.
@@
Beyond addressing the challenge of lacking training datasets whose distribution covers the target solution, another important challenge we often face is the inflexibility of incorporating multi-modal priors into FWI. Along these lines, Wang et al. (2023) demonstrated that injecting priors stored in diffusion models into the inversion process can improve the results. This is done by regularly updating the velocity model using diffusion denoising regularizer after each iteration, leading to increased accuracy of the seismic velocity model. However, the priors used in their approach mainly provide high-level information, reflecting general velocity ranges and structures. While useful, this is often insufficient for achieving the desired level of accuracy in scenarios of poorly acquired data. In this case, additional multi-modal prior information (geological classes, well logs, and structure information) is needed. Hence, the critical challenge lies in how to handle these multi-modal priors flexibly, and that can be addressed by controllable velocity synthesis.
@@
  • We propose to use conditional diffusion models for velocity synthesis.
@@ 2 Methology
Generative models are widely used in seismic exploration, such as in denoising and interpolating seismic data Oliveira et al. (2018, 2019); Gatti and Clouteau (2020); Li et al. (2020); Feng et al. (2021); Wei et al. (2021, 2023), seismic inversion and imaging Mosser et al. (2020); Meng et al. (2021); Zhao et al. (2021); Saraiva et al. (2021); Wang et al. (2023); Yin et al. (2023); Baldassari et al. (2024), interpretation Shi et al. (2020); Durall et al. (2021b, a), and so on. These generative models mainly belong to four types of generative frameworks: variational autoencoders (VAEs) Kingma and Welling (2013), generative adversarial networks (GANs) Goodfellow et al. (2020), normalizing flows Dinh et al. (2014), and diffusion models Sohl-Dickstein et al. (2015). Among all generative models, VAE and normalizing flows require stringent constraints on the model architecture or resort to surrogate objectives for approximating the maximum likelihood training. As mentioned earlier, training GANs is fraught with difficulties. Recently, diffusion models, which circumvent the above limitations, have beaten GANs in generative quality and diversity Dhariwal and Nichol (2021). Here, we choose diffusion models for seismic velocity synthesis. Specifically, to achieve the controllable generation, we use conditional diffusion models.
The fundamental concept behind Denoising Diffusion Probabilistic Models (DDPMs) is straightforward: we aim to acquire the ability to restore images in the 2D context (or videos in the 3D realm) by training denoisers to reverse the corruption process, ultimately restoring the original data distribution from a Gaussian distribution of the latent space. This comprehensive framework can be delineated into two distinct stages: the training phase and the sampling phase, as depicted in Figure 1, where we use seismic velocities as single-channel images. In real-world applications, we aim to exert control over the velocity generation through conditional factors. This control can be introduced during the training process, which is referred to as ‘classifier-free guidance’, or it can be implemented directly in the sampling stage, known as the ‘classifier-guided (reconstruction-guided) diffusion model’. In this section, we will initially delve into the training stage before exploring the sampling process in subsequent subsections.
@@ 2.1 Training process
During the training, given clean data (seismic velocity models) $\boldsymbol{x}\sim p(\boldsymbol{x})$, we use a forward process (Gaussian process) to corrupt the data and obtain the noisy $\boldsymbol{x}_{t}$, $\boldsymbol{x}_{t}\sim p_{t}(\boldsymbol{x})$, which satisfies a linear stochastic differential equation (SDE) (Song et al., 2020b):
@@
where $\beta(t)$ is a differentiable noise scheduler to enforce $\boldsymbol{x}_{1}$ to satisfy a pre-defined noise distribution (here it is the standard Gaussian distribution), $\boldsymbol{w}$ is the Wiener process (Brownian motion, which adds random noise to data at each time step), and $d$ denotes the change in a variable, like $\boldsymbol{x}$, $t$, or $\boldsymbol{w}$ in Equation 1. As shown at the top of Figure 1, for a clean data sample $\boldsymbol{x}_{0}$, we gradually add noise to the data $\boldsymbol{x}_{t}$ until we obtain a final image $\boldsymbol{x}_{1}$ of random noise. The discrete form of Equation 1 (taking DDPM as an example) can be formulated as:
@@
$\mathbb{E}_{\boldsymbol{\epsilon},t}\left[\left(\frac{1}{1-\bar{\alpha}_{i}}\right)^{(-\gamma)}\left\|\Phi\left(\theta,\boldsymbol{x}_{i}\right)-\boldsymbol{x}\right\|_{2}^{2}\right],$ (3)
where $\gamma$ is the loss weight value Choi et al. (2022). In practice, we do not use the direct prediction of $\boldsymbol{x}$ via the neural networks. Instead, we use the $\boldsymbol{\epsilon}$-prediction parameterization, which often results in better sample quality (Ho et al., 2020), defined as $\Phi(\theta,\boldsymbol{x}_{i})=(\frac{\boldsymbol{x}_{i}}{\sqrt{\bar{\alpha}_{i}}}-\sqrt{\frac{1-\bar{\alpha}_{i}}{\bar{\alpha}_{i}}}\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i}))$ and train the $\boldsymbol{\epsilon}_{\theta}$ in $\boldsymbol{\epsilon}$ space. It results in the solution of a scaled score estimate $\boldsymbol{\epsilon}_{\theta}(z_{t})\approx-\sqrt{1-\bar{\alpha}_{t}}\nabla_{\boldsymbol{x}}\log p_{t}\left(\boldsymbol{x}\right)$.
@@ 2.2 Unconditional sampling process
To generate an unconditional sample, a diffusion model starts with a random Gaussian noisy image and then applies a sequence of denoising steps to reduce the noise gradually, as shown at the bottom of Figure 1. At each step, the model predicts a residual image that is added to the noisy image to produce a less noisy image. The process continues until the noise is sufficiently reduced and a high-quality sample is obtained. The sampling process can be described in the continuous form as solving a reverse-time SDE (Song et al., 2020b)
@@ 2.3 Conditional sampling process
In the conditional sampling (conditional generation) setting, the generation process can be formulated in the context that given the conditioning signal $\boldsymbol{c}$, which can represent a class label, smoothness level, well information, imaging results, or any other type of condition. The model generates output $\boldsymbol{x}$ satisfying the distribution of $p(\boldsymbol{x})$ while consistent with the set condition. Two conditional samplers are introduced here, including the reconstruction-guided samplers and classifier-free guidance samplers. The former can be applied to any pre-trained diffusion model without additional training cost, as the guidance only happens at the sampling stage, while the latter requires training the diffusion model to include the conditions.
Reconstruction guided generation. In this way, our target is to make the generated velocity from the unconditional generation sampler $\Phi(\theta,\boldsymbol{x})$ satisfy the conditions, in which we can use an objective function to guide the process at each time step $i$:
@@
where $mask$ denotes the mask operation on the generated samples to extract, for our specific example, the well location information to compare with the condition given by the true well log. We are using this objective function to calculate the gradient with respect to $\boldsymbol{x_{i}}$ to update the sampler $\Phi(\theta,x_{i})$. Then, the final samples at each time step $t$ is
$\hat{\Phi}(\theta,\boldsymbol{x_{i}})\leftarrow\Phi(\theta,\boldsymbol{x}_{i})-\frac{w_{r}\sqrt{\bar{\alpha}_{i}}}{2}\nabla_{\boldsymbol{x}_{i}}\left\|\boldsymbol{c}-mask(\Phi\left(\theta,\boldsymbol{x}_{i}\right))\right\|_{2}^{2},$ (7)
where the additional term with reweighting in this equation compared to the original sampler $\Phi(\theta,\boldsymbol{x}_{i})$ is the guidance based on the model’s reconstruction of the condition, so-called reconstruction-guided sampling Ho et al. (2022), $w_{r}$ is the weighting factor to control the strength of the condition, the left arrow to denote the replacement of the original unconditional sampling procedure $\Phi(\theta,\boldsymbol{x}_{i})$ by the conditional sampling implementation $\hat{\Phi(\theta,\boldsymbol{x}_{i})}$. With this implementation, the trajectory of the reverse process is modified by the gradient of the objective function to guide the generated velocity to satisfy the prior distribution as well as the given well condition. However, this process is very sensitive to the choice of the weighting factor, which we will demonstrate via the numerical results, yielding inflexible control.
Classifier-free guidance sampler. When considering the classes of the images, a concept similar to the reconstruction-guided diffusion model is called classifier guidance diffusion, whose generation process is guided by the gradient from an external classifier. Although it can feasibly guide the unconditional diffusion model with a plug-in classifier, this process will increase the computational cost as it requires gradient calculation during the inference. Another drawback of this type of method is that the generation sometimes will ignore the guidance of the classifier via the adversarial noise, and then the generation, in this case, will actually be unconditional. Another controllable sampling is classifier-free guidance sampling. Compared to reconstruction-guided sampling, this requires training the diffusion model with the condition as part of the training. The good news is that it often has higher-quality generation during the sampling stage. Our training target is to train a neural network to fit the distribution $p(\boldsymbol{x}|\boldsymbol{c})$ given a conditioning signal $\boldsymbol{c}$. Compared to the previously mentioned diffusion model (Figure 1), the only modification is changing the $\Phi(\theta,\boldsymbol{x})$ to $\Phi(\theta,\boldsymbol{x},\boldsymbol{c})$. To improve the sample quality, instead of directly using the prediction $\Phi(\theta,\boldsymbol{x},\boldsymbol{c})$, we can use the following weighting form with the ability to adjust the strength of the conditioning,
@@
where $\lambda$ is the conditioning scale, and $\Phi(\theta,\boldsymbol{x}_{i},\boldsymbol{c})=(\frac{\boldsymbol{x}_{i}}{\sqrt{\bar{\alpha}_{i}}}-\sqrt{\frac{1-\bar{\alpha}_{i}}{\bar{\alpha}_{i}}}\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i},\boldsymbol{c}))$ is the conditional model prediction. When $\lambda$ is positive, the model output will have lower diversity, but higher quality (high-fidelity) compared to the case where the output is only dependent on the $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i},\boldsymbol{c})$ (Ho and Salimans, 2022). To make the diffusion model more flexible, in training, we can randomly add embedding vectors of the condition as well as the vectors of a null condition denoting unconditioning, yielding flexible generation abilities when sampling. With this trick in the training stage, the unconditioned $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i})$ can be realized by setting $\boldsymbol{c}$ as a vector of null condition in $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{i},\boldsymbol{c})$. Next, we will discuss how to incorporate the condition in the neural networks.
Figure 2: Conditional generative diffusion models. The backbone network is U-net (Ronneberger et al., 2015).
As shown in Figure 2, there are two ways to achieve the classifier-free guidance. In the first way, the class label or the well log information is embedded as a vector and then added to the hidden layer of the U-Net. Specifically, for the class labels, we start an embedding layer to convert the class to a vector, followed by a Multilayer perceptron (MLP) to obtain an embedded vector while for the well log, we use an MLP to embed the well logs and sinusoidal positional encoding to embed the well location, yielding the embedded vectors. However, these MLPs have limits in handling high-dimensional conditions, e.g., multiple wells or seismic images. In the second way, for 2D input data for conditioning, we utilize cross-attention to add conditional information into the neural networks for training. The concept is based on calculating the similarity between the conditioning data and the hidden output of the U-Net and then updating the latent vectors in the U-Net. We map the tokenizations (specifically for seismic velocity, it is a way to transform the image to the embedded vectors) of the original conditions $\mathbf{c}$, $\tau_{\theta}(\mathbf{c})$, to the intermediate output of the U-Net by implementing cross-attention$(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d}})\cdot V$, between the U-Net features and the input condition (Rombach et al., 2022), where
@@
Here, $\phi_{j}(\boldsymbol{x}_{i})$ is the output of the intermediate layer of the U-Net, $\tau_{\theta}$ used in this paper is $1\times 1$ convolutional layer, and $W_{j}^{Q}$, $W_{j}^{K}$, and $W_{j}^{V}$ are the learnable weight matrices for the query, key, and value, respectively. The cross-attention mechanism enables the model to consider and incorporate information from given conditions while processing the velocity synthesis. To achieve an accurate control of velocity synthesis, we add the spatial information of the velocity (the coordinates of the grid points) as additional channels and then concatenate them with conditions $\boldsymbol{c}$ to feed into the $\tau_{\theta}$. The beauty of this method is that it can incorporate multi-model conditioning. What’s more, thanks to the cross-attention, we can achieve pixel-level controllable generation.